Determine the point estimate of the population​ proportion, the margin of error for the following confidence​ interval, and the number of individuals in the sample with the specified​ characteristic, x, for the sample size provided. Lower bound =0.516​, upper bound =0.834​, n=1000
Answer #1

Solution :

Point estimate = = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = = (0.516+0.834) / 2=0.675

Margin of error = E = Upper confidence interval - = 0.834-0.675=0.159

The number of individuals in the sample with the specified characteristic is ,

x = * n = 0.675*1000

The number of individuals = 675